# InteLiPlan: An Interactive Lightweight LLM-Based Planner for Domestic Robot Autonomy

### v2

@@ IV-C Evaluation of Robustness in Failure Recovery
Figure. 5 demonstrate our framework’s performance in replanning in various scenarios. While ‘task confusion’ mainly processes user preference (e.g. user prefers the robot to get a coke to a 7up), ‘planning failure’ and ‘execution failure’ cases assess the ability of mapping language to the list of pre-defined motion (e.g. output ‘open cupboard’ in response to ‘the object is in the cupboard’). The lower success rates in recovering from ‘execution failure’ are due to the longer sequence of recovery than what was provided in the fine-tuning data. In general, prompting methods fails to embody LLM with sufficient robot-domain understanding. Our method significantly increases the performance of LLM models in robotic task failure recovery, indicating robustness in real-world conditions where adaptability and responsiveness to failures are crucial for operational success.
@@ IV-D Evaluation of Scalability
In this experiment, we evaluate the scalability of the approach by the 101 task instructions from SayCan [1]. To accommodate the task variants, we fine-tuned Mistral with the tasks outlined in Table III, which covers the expected actions required for the dataset solutions.
@@
IV-E Real-robot experiments
We validate the efficiency of the system on the physical HSR. We implemented 3 sets of experiments:
@@ IV-E Real-robot experiments
  • Multi-step failure task: the robot must detect failures and replan twice to be able to complete the task. Specifically, we implement the planning failure from vision first, then after the targeted object is found, it is out of reach and the robot should ask again to complete the task.
The result demonstrates that by processing both vision and reachability capabilities, all planned action sequences can be executed with collision-free trajectories. Fig. 1 and Fig. 6 showcase InteLiPlan results in real-world settings, which includes tasks with and without failure recovery. The multi-step failure verifies that our method enables autonomy with seamless end-user intervention.
@@
Table V records the averaged planning time of the system over the trials, which is divided into vision, feasibility and planner queries processing time. In total, the model can plan an executable action sequence in less than 7s. The feasibility check module, which takes up the longest processing time, can be relaxed with fewer nodes in the reachability graph. Notably, using YOLO instead of VLM for vision input significantly contributes to real-time and onboard processing capability of the pipeline.
Videos of our demonstrations are available at our project page: https://kimtienly.github.io/InteLiPlan.
TABLE V: Breakdown of onboard planning time (s).
Vision query Feasibility query Planner query
6e-6 5 1.5
@@ V Conclusion
We presented InteLiPlan, an interactive lightweight LLM-based robotics planner for reliable and robust autonomy in domestic environments. Our framework employs a conversation-like format between internal modules and facilitates human-robot interaction, allowing the system to reason about failures through a multimodal input formulation. By incorporating a human in the loop, the robot gains the ability to replan based on instructions, effectively utilizing human input for dynamic adjustments. We performed extensive evaluations to investigate common failures that happen in real scenarios, where human instruction is required. The interactive behaviors guarantee interpretability and reliability for robots operating in domestic environments.
With roughly 300 data samples for fine-tuning, our approach effectively handles the targeted tasks by leveraging the human-like text understanding capabilities of LLMs. This reduces the effort of defining traditional task planning domains in robotics, provides an intuitive, robot-independent data structure and boosts applicability in resource-constrainted applications. Notably, deploying our approach with the lightweight Mistral 7B model achieves both comparable results with the SOTA baseline and –notably to our approach– real-time onboard computing. Our system was validated on the physical Toyota HSR robot.

### v3

@@ IV-A3 Evaluation of Robustness in Failure Recovery
Fig. 5 demonstrates our framework’s performance in replanning in various scenarios. While ‘task confusion’ mainly processes user preference (e.g. user prefers the robot to get a coke to a 7up), ‘planning failure’ and ‘execution failure’ cases assess the ability of mapping language to the list of pre-defined motions (e.g. output ‘open cupboard’ in response to ‘the object is in the cupboard’). The lower success rates in recovering from ‘execution failure’ are due to the longer sequence of recovery than what was provided in the fine-tuning data. In general, prompting methods fails to embody LLM with sufficient robot-domain understanding. Our method significantly increases the performance of LLM models in robotic task failure recovery, indicating robustness in real-world conditions where adaptability and responsiveness to failures are crucial for operational success.
The w/o failure column of Table II reports success rates in scenarios where the planner must succeed in the two-step process of failure detection followed by failure recovery. The results reveal a significantly increased gap between prompt-based models and our approach as opposed to w/o failure cases, underscoring the effectiveness of our method in handling multi-step interaction and ensuring task completion.
@@
IV-B Evaluation of Scalability Across Tasks
In this experiment, we evaluate the scalability of the approach by 101 task instructions from SayCan [6]. To accommodate the task variants, we fine-tuned Mistral with the tasks outlined in Table III, which covers the expected actions required for the dataset solutions.
This set of experiments added a metric of Execution success rate, which is the ratio of the trials where the robot successfully executes the planned actions. Given that SayCan [6] does not have failure recovery ability, we only report the result of cases without failure recovery. Generative result from PaLM 8B [38] is also reported as a lightweight LLM baseline.
@@ IV-B Evaluation of Scalability Across Tasks
IV-C Evaluation of Cross-embodiedment with Real-robot Experiments
We validate the cross-embodiment feature of the system on two different physical robot platforms that are different in both camera setups and the number of degrees of freedom (DoF). In both robot setups, InteLiPlan is processed entirely onboard with a single NVIDIA Geforce RTX 2070, verifying onboard computing capability. We tested a pre-trained InteLiPlan model on both platforms:
  • Toyota Human Support Robot (HSR): 11-DoF robot. The vision module takes the signal from the Asus Xtion PRO Live stereo camera on the head of the robot.
  • Anymal D-Unitree Z1 (Anytree): 12-DoF robot (where the Anymal D is simplified with a planar base). The vision module takes the signal from the Realsense D435 on the wrist of the end effector.
We implemented 3 sets of experiments, with 10 trials each:
@@ IV-C Evaluation of Cross-embodiedment with Real-robot Experiments
  • Multi-step failure task: the robot must detect failures and replan twice to be able to complete the task. Specifically, we implement the failure from vision first, then after the targeted object is found, it is out of reach and the robot should ask again to complete the task.
TABLE V: Success Rates (%) of Task Planning.
Robot w/o failure 1-step failure Multi-step failure
HSR 100 100 90
Anytree 100 100 90
Table V recorded the success rates of InteLiPlan in generating the correct sequences of action upon the corresponding case. The success rates were also affected by the success of vision and feasibility modules, leading to the failed cases. The result demonstrates that with one model, planned action sequences guarantee executability across embodiments. Fig. 1 and Fig. 6 showcase InteLiPlan results in real-world settings, which include tasks with and without failure recovery. The multi-step failure verifies that our method enables autonomy with seamless end-user intervention.
@@
Table VI records the averaged planning time of the system over the trials, which is divided into vision, feasibility and planner queries processing time. In total, the model can plan an executable action sequence in less than 7s with onboard resources. The feasibility check module, which takes up the longest processing time, can be relaxed with fewer nodes in the reachability graph. Notably, using YOLO instead of VLM for vision input significantly contributes to real-time and onboard processing capability of the pipeline.
TABLE VI: Breakdown of average onboard planning time(s).
Robot Vision query Feasibility query Planner query
HSR 6e-6 5 1.5
Anytree 8e-6 5.5 1.3
The experiment validates the cross-embodiedment feature that requires no extra finetuning, further strengthening that InteLiPlan’s planning capability is universal to different robots, and also different camera views, as opposed to vision-language-action models [39].
@@ V Conclusion
We presented InteLiPlan, an interactive lightweight LLM-based robotics planner for reliable and robust autonomy in domestic environments. Our framework employs a conversation-like format between internal modules and facilitates human-robot interaction, allowing the system to reason about failures through a multimodal input formulation. By incorporating a human in the loop, with human-like text understanding capabilities of LLMs, the robot gains the ability to replan based on instructions, effectively utilizing human input for dynamic adjustments. We performed extensive evaluations to investigate common failures that happen in real scenarios, where human instruction is required. The interactive behaviors guarantee interpretability and reliability for robots operating in domestic environments.
This paper reduces the effort of defining traditional task planning domains in robotics, provides an intuitive, robot-independent data structure and boosts applicability in resource-constrainted applications. Notably, deploying our approach with the lightweight Mistral 7B model achieves both comparable results with the SOTA baseline and –notably to our approach– real-time onboard computing. Our system was validated on both the physical HSR and Anymal D-Unitree Z1 robots, verifying the cross-embodiedment capability across different DoFs and camera views.